THE MILLIMETER AND SUBMILLIMETER-WAVE ROTATIONAL SPECTRUM OF THE 8^{1}9^{1}, 6^{1}7^{1}, AND 2^{1} VIBRATIONAL STATES OF NITRIC ACID

\underline{DOUGLAS T. PETKIE}, Department of Physics, Wright State University, Dayton OH 45435; PAUL HELMINGER, Department of Physics, University of South Alabama, Mobile, AL 36688; IVAN MEDVEDEV, ATSUKO MAEDA, and FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210.

The measurement and analysis of the rotational spectrum of nitric acid in 8^{1}9^{1}, 6^{1}7^{1}, and 2^{1} vibrational states will be presented. Measurements over the frequency range of 78-118 GHz were made with a phased locked BWO synthesizer and measurements from 118-520 GHz were made using the FASSST spectrometer with a heated cell. The rotational transitions in the 8^{1}9^{1} and 6^{1}7^{1} states exhibit a ∼3.5 MHz and ∼1.5 MHz torsional splitting, respectively, and a set of torsional parameters were required to fit the spectra. A Fermi interaction between the 8^{1}9^{1} and 6^{1}7^{1} states was also included in order to fit all of the observed transitions. There is no resolvable torsional splitting in the 2^{1} state and it is free from strong perturbations within the range of measured transitions. The assignments and on-going analyses of several other vibrational states will also be presented.